OpenStudy (anonymous):

Find the first,fourth,and tenth terms of the arithmetic sequence described by the given rule. A(n) = 12+(n-1)(3)

OpenStudy (whpalmer4):

\[a_n = 12+(n-1)(3)\]The first term is \(a_1\) and \(n = 1\): \[a_1 = 12 + (1-1)(3) = 12 + 0(3) = 12\] You should be able to find \(a_4, a_{10}\) which have values of \(n = 4, n=10\) respectively (I hope you see the relationship!)

OpenStudy (anonymous):

Thank you

OpenStudy (whpalmer4):

Do you have some answers for me to check?

OpenStudy (anonymous):

Yes but this is a different question if a snail travels at a rate of 2.37 feet per minute wouldn't the rule be 2.37×6

OpenStudy (anonymous):

And it ask how far will of travel in 6 minutes I put 14.02

OpenStudy (anonymous):

14.22 I meant